# High-Precision Transition Energy Measurements of Neonlike Fe XVII Ions

### v2

@@ 3 Discussion of the results
aThis column shows the difference of previous theoretical large CI computations from Kühn et al. (2022) with present experiment.
@@
bThis column shows the difference of current theoretical large CI computations with present experiment.
@@
We compare the present results in Fig. 2 and Tabs. 1, 2 with earlier experimental data, observations, and predictions, including our own. Our calculations employ the latest version of our highly scalable parallel CI code (Cheung et al., 2021) (see Appendix). Optimization of the basis set construction allowed faster convergence with the principal quantum number than in our prior work (c. f. supplementary material of Kühn et al. (2022)), while including higher partial waves (, , and ), and larger number of reference configurations of even and odd parity. Table 3 shows the QED and other contributions in cm-1 for the measured transitions. Column shows the results obtained with the basis set (see the Appendix); column , additional contributions from highly excited orbitals up to ; “Extras”, additional contributions due to the much larger number of configurations included in CI, selected to give the large contributions. The final results are the sum of these three columns and a QED contribution (Tupitsyn et al., 2016). Columns [This Th.b] show differences between the current experiment and theory, and [Prev. Th.a] show differences from previous calculations presented in (Kühn et al., 2022), demonstrating a significant improvement. We estimate the uncertainty in the electronic correlation computations to be approximately 29 meV (230 cm-1), primarily arising from the contribution (see Appendix). The difference between theory and experiment is within the combined uncertainties for all six levels. This allows us to estimate the uncertainty of the QED contribution at 30-32 meV (240 - 260 cm-1), which is the combined theory and experimental uncertainties added in quadrature. We also computed the energies of the 3A and 3B levels for the first time. The lines (3A and 3B), which involve a electron, have the largest QED contributions ( meV, or cm-1), while for transitions 3G and 3F as well as ones (3C and 3D) they are much smaller. From the 3A and 3B results, we thus estimate a relative QED accuracy of 8%.
@@
As shown in Tab. 1, line 3F, close to the He-like F K calibration line, shows a larger absolute deviation of about 33 meV from the large CI prediction, while the remaining measured lines remain below 10–20 meV. Unfortunately, line 3F was measured only once, unlike the others, which were scanned at least 4–5 times. We explored several plausible explanations for the 3F discrepancy. One possible source of the discrepancy could be the simultaneous excitation of high Rydberg lines of O VII within the scan range of line 3F, which could lead to a shift of the 3F centroid. Furthermore, we considered lines from the lower charge states, Fe X, Fe VIII, and Fe VII, which fall within the 3F scan range. Despite the relatively low abundance of these charge states in our experiment, they can potentially influence the 3F line due to their strong oscillator strengths. Although the theoretical line positions and oscillator strengths of these low charge states are calculated by (Gu et al., 2006), they have never been compared experimentally, making it difficult to estimate their influence on the 3F position. We also investigated the possibility of magnetically induced mixing of the and excited states (Beiersdorfer et al., 2003), which might shift the energy of the state sufficiently to introduce a systematic error in our measurement of 3F. However, measurements by Beiersdorfer et al. (2016) show a separation of 1.2 eV between these states, making strong magnetic-field-induced mixing unlikely. We performed FAC calculations for atoms in strong magnetic fields to verify this, finding shifts on the order of 10 µeV for the field strength in PolarX-EBIT, demonstrating that this effect is not important in our experiment. The decrease in reflectivity of the platinum-coated diffraction grating over the 3F scan range could slightly affect the centroid position determination at 739 eV. Based on simulations we estimate this effect to be smaller than 0.1 meV.
@@
We also consider the differences of the three line pairs, as they are more sensitive to QED effects than absolute energies. Table 3 shows that the largest uncertainty, caused by the uncertainty in the basis-set convergence, is common to each of the pairs. This significantly reduces the uncertainty of electronic correlations to better than 6 meV (50 cm-1) for the energy difference. Both (3A-3B) and (3C-3D) are in excellent agreement with our present as well as previous predictions (Kühn et al., 2022). For (3G-3F), the deviation is 46 meV (about ) and can be attributed to the factors discussed above for the line 3F. It is interesting to note that our measured 3F energy is in much better agreement with solar observations (Phillips et al., 1982; Del Zanna & Ishikawa, 2009) than with our calculations. Nevertheless, our present calculations of the ground state transitions show an order of magnitude smaller deviation from our experimental results compared to our prior predictions (Kühn et al., 2022). This represents a benchmark with our experimental data at the level of 10–20 ppm, an unprecedented agreement for a neonlike system to the best of our knowledge.
@@
Besides CI, we performed calculations using a combination of conventional CI and second-order many-body perturbation theory (MBPT) with the Flexible Atomic Code (FAC) (Gu, 2008). Details of this method are described in Gu (2005); Gu et al. (2006), and recently in Steinbrügge et al. (2022). In these calculations, we included frequency-dependent generalized Breit interactions in both the CI expansion and the MBPT corrections, as well as self-energy and vacuum polarization calculated using the QED operator model of Shabaev et al. (2018). These predictions demonstrate a reasonable agreement with our experimental data, with the largest discrepancy of about 100 meV observed for line 3C. We compared our results with other CI+MBPT data available in the literature (Wang et al., 2016; Santana et al., 2015; Gu, 2005), showing maximum deviations of up to 0.5 eV. The origin of the discrepancy between our CI+MBPT calculations and the previously published ones is unclear. We have also observed departures from the predictions of multi-configuration Dirac-Fock (MCDF) and autostructure (AS) calculations (Aggarwal et al., 2003; Loch et al., 2005; Liang & Badnell, 2010), with deviations reaching up to 1–3 eV. However, we note that the atomic structures used in these calculations were necessarily small to facilitate their use in R-matrix collision calculations, which are computationally more demanding compared to atomic structure calculations. Other accurate predictions for Fe XVII from multi-reference Møller-Plesset (MRMP) are reported in Del Zanna & Ishikawa (2009) and included in the CHIANTI code. They show very good agreement with our experimental data.
@@
We compare our results with laboratory data from Beiersdorfer & Wargelin (1994) and Brown et al. (1998). Both works measured electron impact spectra of Fe XVII under similar experimental conditions in the Lawrence Livermore National Laboratory (LLNL) EBIT facility using a crystal spectrometer employing a cesium acid phthalate (CsAP) crystal for the wavelength range of the lines presently discussed. Both measurements have carefully concatenated several spectra from different wavelength ranges and calibrated them against reference lines of hydrogenic and heliumlike oxygen, fluorine, and neon, similar to our work. Nevertheless, these two measurements are themselves marginally inconsistent with each other within their quoted uncertainties. Furthermore, we find that our measurements are also marginally inconsistent with both these previous measurements within uncertainties. The source of these marginal inconsistencies is unknown.
@@
We also compare our results with data from widely used databases and plasma codes. The NIST Atomic Spectroscopy Database (ASD) (Kramida et al., 2022) values showed significant deviations for lines 3C and 3A. However, when critically evaluated data by the authors of the NIST ASD (Kramida, 2019) were considered, we found a much better agreement with our experimental results, see Tab. 2. Comparison with AtomDB (Foster et al., 2012), CHIANTI (Del Zanna et al., 2021), and SPEX (Kaastra et al., 1996) databases and plasma codes revealed discrepancies as large as 1–2 eV. SPEX numbers showed better agreement with our results than those found in AtomDB, since SPEX has updated Fe-L atomic data (Gu et al., 2019, 2020, 2022), which were mainly calculated using FAC. Although the Astrophysical Plasma Emission Database (APED) version in AtomDB shows different values in its online webguide version (2.0.1) 222http://www.atomdb.org/Webguide/webguide.php and its pyatomdb version (3.0.4), the theoretical source is in both cases Loch et al. (2005), which uses the CI method, and disagrees by up to 1–2 eV from our results, as shown in Tab. 2. CHIANTI provides two sources for the energies: Autostructure (AS) theory (Liang & Badnell, 2010), and the more accurate set of data from MRMP theory (Del Zanna & Ishikawa, 2009). Note that the theoretical energy level data in AtomDB and Chianti are not used when generating model spectra, when more accurate experimental or observational values exist. AtomDB replaces the most important transition energies of Fe XVII with the Brown et al. (1998) values, whereas CHIANTI uses the observed transition energies from solar Observation (Del Zanna & Ishikawa, 2009).
@@
Overall, most experimental and observational data agree with our experiment within 0.1 eV on average, well within error bars of earlier works. However, there are substantial discrepancies with predictions from certain theoretical models, exceeding the margins of error associated with the experimental results. This highlights the urgent need to update the aforementioned databases to avoid pitfalls in astrophysical spectrum modeling and interpretation of observational data.
@@ 4 Summary and Conclusions
We presented high-precision transition-energy measurements of eight strong, astrophysically preeminent Fe XVII transitions required for plasma diagnostics. Our approach combined resonant photoexcitation of Fe XVII and narrow H-like and He-like transitions with high-resolution photoelectron spectroscopy (Rossnagel et al., 2001). This eliminates a very common source of systematic errors found even in advanced monochromators, namely quasi-periodic encoder interpolation errors (Follath & Balzer, 2010; Krempaský et al., 2011). As a result, our Fe XVII measurements represent a significant improvement in accuracy compared to previous experimental references, achieving an average enhancement of almost an order of magnitude. The uncertainties now stand at 10–15 meV, which translates to Doppler shifts of approximately km/s. A further improvement in accuracy by another order of magnitude will require incorporating high-resolution/-stability voltage sources and more accurate voltmeters at ASPHERE to eliminate systematic errors associated with knowledge of the bias voltages.
@@
We have also improved our high-precision calculations by an order of magnitude in comparison with previous best calculations (Kühn et al., 2022; Cheung et al., 2021). This improvement allowed us, for the first time, to test the accuracy of QED corrections to the transition energies of a complicated 10-electron system. We expect that the achieved QED accuracy is applicable to a broad range of ions of intermediate degrees of ionization that can be treated with our large-scale CI or CI+all-order approaches. This has significant implications for predicting energy levels in systems where no experimental data are available for a wide range of applications in astrophysics, plasma physics, and atomic clocks development (King et al., 2022). The established QED accuracy is deemed sufficient for high-precision prediction of HCI clock transitions (Kozlov et al., 2018). Improved accuracy of the experimental values would allow to further decouple uncertainty due to basis-set convergence from the uncertainty in the QED and improve theory predictions.
@@
Our improved transition energy measurements for Fe XVII are sufficiently accurate that the uncertainties are no longer a significant part of the error budget for present or future planned astrophysical instruments, such as Chandra HETGS, XMM-Newton RGS, XRISM (Tashiro et al., 2018), Athena (Pajot et al., 2018), LEM (Kraft et al., 2022), HUBS (Cui et al., 2020), Arcus (Heilmann et al., 2022), HiReX (Nicastro et al., 2021), and Lynx (Schwartz et al., 2019). Future campaigns of similar measurements of prominent transitions in key ions (especially Fe L-shell ions) would be of great utility and could easily be directly included in commonly used astrophysical plasma spectral databases.
@@
The closeness of our large CI calculations to our measured values, with the worst deviation at line 3F of 33 meV amounting to a Doppler shift of only 13 km/s, shows that such well-converged calculations are sufficiently accurate to be readily used in spectral databases. While there is no reason for this in the case of the lines measured in the present work, when accurate measurements are not available, similarly well-converged results could be used for other transitions of Fe XVII and many other ions. By including a very large number of configurations, our agreement becomes significantly better than that of other well-performing methods, such as results from less-converged large CI, MBPT and MRMP calculations. For Fe XVII, our calculations are more accurate than even the best measurements for Fe L-shell transitions in Li-like through F-like ions (Brown et al., 2002). This suggests a near-future research program composed of comprehensive large-CI calculations of transition energies for all ions of astrophysical interest up to Ne-like, supplemented by targeted experiments aimed at measuring the most important transition energies.
@@ Appendix A Large-scale CI calculations
The low-lying energies and wave functions are determined by solving the time-independent multi-electron Schrödinger equation
Expanding the previous work (Kühn et al., 2022), we perform several calculations optimizing the basis set convergence, including higher partial waves up to orbitals, and significantly expanding the set of reference configurations until convergence is reached in these parameters as well.
We have shown a comparison of our theoretical results for six transitions measured in this work with the experiment in the main text. We note that such a larger scale computation of the 4C and 4D levels is beyond the capabilities of available computational resources (32 TB of memory and about 2000 CPUs on our largest available partition). Computing higher-lying levels requires computing of all the lower lying levels with the same angular momentum and parity, drastically increasing memory requirement.
@@
We will discuss a complete assessment of the main contributions to the energies, including the basis set construction, the inclusion of extra configurations, and QED. We find excellent agreement with experiments for all energies, at the level of 0.0004% for some levels. With the high level of accuracy attained, we are able to test QED contributions in the calculations of multi-electron systems for the first time.
@@
The base calculation for the energy levels is done with a basis set and is listed in cm-1 in Tab. 4 in column “”. The contributions to the energy levels from expanding the basis set to and are in the columns “” and “”, respectively. The largest difference between the and calculations was 3 cm-1, so the basis set at the level of partial waves is considered sufficiently saturated. We note that although the basis was also used in Kühn et al. (2022), we constructed more compact basis in the present work, to significantly improve convergences with the principal quantum number . The basis is constructed in the 5 a.u. cavity, while the basis in Kühn et al. (2022) was constructed in a 20 a.u. cavity, with additional differences in the constructions of the higher partial wave orbitals. Detailed comparison of the two computations confirms much better convergence properties of the present basis. We note very large computational resources needed for a basis set expansion, especially for the inclusion of higher partial waves.
Contributions to higher partial waves are considered in the next six columns of Tab. 4. We calculated the contributions of extending the base basis set to include up to orbitals and list them in column “”. Next, we successively increase the principal quantum number and increase the basis set up to . The contributions from orbitals and orbitals are given in columns “” and “”, respectively. The largest difference between and calculations was 9 cm-1, so the energies of including the higher orbitals have also converged sufficiently. We note that a large fraction of the contribution comes from very high orbitals, so inclusion of the first few orbitals does not give correct results for this partial wave. This effect is exacerbated for the and orbitals, where more of the contribution is expected to come from even with the present compact basis.
@@
We note that we have performed detailed convergence studies computing a separate contribution for each for the last few principal quantum numbers to evaluate convergence. Based on these data, we conservatively estimate the missing higher orbital contribution at 5 cm-1, higher orbital contribution at 20 cm-1, and higher orbital contribution at 50 cm-1. It appears that is not sufficiently converged. Table 4 shows that the contribution of all orbitals is about 1/2 of the contribution. Conservatively assuming a similar convergence pattern for higher partial waves gives 70 cm-1 for the partial wave and a similar total contribution for all the other partial waves. The total uncertainty due to the convergence of the basis set is then of the order of 230 cm-1. However, we note that the incomplete convergence of the basis is expected to cause a systematic shift of data for all levels. i.e. all energy values will be larger, with some smaller variances between the levels. It is possible that the partial wave convergence is faster and the overall shift is smaller; therefore, we only use the above estimate to make an accuracy evaluation, but do not shift the theory values. We note that overall +100 cm-1 shift of all of our values would improve the agreement of our data with the present experiment, however, this is the level of the experimental precision at 1 so improved experimental precision is needed to definitively test the basis set convergence.
@@
Additionally, an extensive evaluation of the configuration weights was done to include important configurations in the list of basic reference configurations used to construct a final set of configurations. The weights of configurations signify the configuration’s contribution to the corresponding wave functions, and are calculated for each configuration as from Equation A1. These calculations are done by allowing single and double excitations to a much smaller basis set, since the size of the computational problem will become prohibitive when additional reference configurations are included. The total contributions to including these extra configurations are given in the column “Extras” in Tab. 4. Beyond the initial 2 even- and 3 odd-parity configurations, we systematically included an additional 12 even- and 9 odd-parity reference configurations. Note that energies were calculated only for 2 even-parity levels to save computational resources and allow additional reference configurations. The inclusion of these extra configurations contributes about 100 cm-1 shift to the energies and accounts for an additional 2 million relativistic configurations. Note also that these contributions would also be higher if the calculations were done with a larger basis set. We estimate an uncertainty from the convergence of the CI configuration set at the level of 50 cm-1, which is essentially negligible in comparison with the basis set convergence uncertainty. We note that missing contributions can be both positive and negative in this case.
@@
Analysis of contributions to the 3F-3G, 3C-3D, and 3A-3B line differences given in main text Table 3 shows that the basis set expansion contribution effectively cancels for similar configurations; it is less than 50 cm-1 for all three cases. We also find that QED contributions play a major role in the 3F-3G energy difference. For 3C-3D, the contributions from the basis set expansion and addition of extra configurations essentially cancel out the QED. In the 3A-3B difference, the basis set and extra configuration contributions cancel, leaving a shift from the QED. Therefore, comparing the differences of the energy values for similar configurations provides important additional information. It would be very useful to improve the uncertainty of the experiment as well as carry out such comparison in other ions with different degrees of ionization with 7-10 electrons.

### v3

@@ 3 Discussion of the results
aThis column shows the difference between previous theoretical large CI computations from Kühn et al. (2022) and those of the present experiment.
@@
bThis column shows the difference between current theoretical large CI computations and those of the present experiment.
@@
We compare the present results in Fig. 2 and Tabs. 1 and 2 with earlier experimental data, observations, and predictions, including our own. Our calculations employ the latest version of our highly scalable parallel CI code (Cheung et al., 2021) (see Appendix). Optimization of the basis-set construction allowed faster convergence with the principal quantum number than in our prior work (see supplementary material of Kühn et al. (2022)), while including higher partial waves (, , and ), and a larger number of reference configurations of even and odd parity. Table 3 shows the QED and other contributions in cm-1 for the measured transitions. Column shows the results obtained with the basis set (see the Appendix); column , additional contributions from highly excited orbitals up to ; “Extras”, additional contributions due to the much larger number of configurations included in CI, selected to give the large contributions. The final results are the sum of these three columns and a QED contribution (Tupitsyn et al., 2016). Columns [This Th.b] show differences between the current experiment and theory, and [Prev. Th.a] show differences from previous calculations presented in Kühn et al. (2022), demonstrating a significant improvement. We estimate the uncertainty in the electronic correlation computations to be approximately 29 meV (230 cm-1), primarily arising from the contribution (see Appendix). The difference between theory and experiment is within the combined uncertainties for all six levels. This allows us to estimate the uncertainty of the QED contribution at 30–32 meV (240–260 cm-1), which is the combined theory and experimental uncertainties added in quadrature. We also computed the energies of the 3A and 3B levels for the first time. The lines (3A and 3B), which involve a electron, have the largest QED contributions ( meV, or cm-1), while for transitions 3G and 3F as well as ones (3C and 3D) they are much smaller. From the 3A and 3B results, we thus estimate a relative QED accuracy of 8%.
@@
As shown in Tab. 1, line 3F, close to the He-like F K calibration line, shows a larger absolute deviation of about 33 meV from the large CI prediction, while the remaining measured lines remain below 10–20 meV. Unfortunately, line 3F was measured only once, unlike the others, which were scanned at least 4–5 times. We explored several plausible explanations for the 3F discrepancy. One possible source of the discrepancy could be the simultaneous excitation of high- Rydberg lines of O VII within the scan range of line 3F, which could lead to a shift of the 3F centroid. Furthermore, we considered lines from the lower charge states, Fe X, Fe VIII, and Fe VII, which fall within the 3F scan range. Despite the relatively low abundance of these charge states in our experiment, they can potentially influence the 3F line due to their strong oscillator strengths. Although the theoretical line positions and oscillator strengths of these low charge states are calculated by Gu et al. (2006), they have never been compared experimentally, making it difficult to estimate their influence on the 3F position. We also investigated the possibility of magnetically induced mixing of the and excited states (Beiersdorfer et al., 2003), which might shift the energy of the state sufficiently to introduce a systematic error in our measurement of 3F. However, measurements by Beiersdorfer et al. (2016) show a separation of 1.2 eV between these states, making strong magnetic-field-induced mixing unlikely. We performed FAC calculations for atoms in strong magnetic fields to verify this, finding shifts on the order of 10 µeV for the field strength in PolarX-EBIT, demonstrating that this effect is not important in our experiment. The decrease in reflectivity of the platinum-coated diffraction grating over the 3F scan range could slightly affect the centroid position determination at 739 eV. Based on simulations we estimate this effect to be smaller than 0.1 meV.
@@
We also consider the differences between the three line pairs, as they are more sensitive to QED effects than absolute energies. Table 3 shows that the largest uncertainty, caused by the uncertainty in the basis-set convergence, is common to each of the pairs. This significantly reduces the uncertainty of electronic correlations to better than 6 meV (50 cm-1) for the energy difference. Both (3A–3B) and (3C–3D) are in excellent agreement with our present as well as previous predictions (Kühn et al., 2022). For (3G–3F), the deviation is 46 meV (about ) and can be attributed to the factors discussed above for the line 3F. It is interesting to note that our measured 3F energy is in much better agreement with solar observations (Phillips et al., 1982; Del Zanna & Ishikawa, 2009) than with our calculations. Nevertheless, our present calculations of the ground state transitions show an order of magnitude smaller deviation from our experimental results compared to our prior predictions (Kühn et al., 2022). This represents a benchmark with our experimental data at the level of 10–20 ppm, an unprecedented agreement for a neon-like system to the best of our knowledge.
@@
Besides CI, we performed calculations using a combination of conventional CI and second-order many-body perturbation theory (MBPT) with the Flexible Atomic Code (FAC) (Gu, 2008). Details of this method are described in Gu (2005); Gu et al. (2006), and recently in Steinbrügge et al. (2022). In these calculations, we included frequency-dependent generalized Breit interactions in both the CI expansion and the MBPT corrections, as well as self-energy and vacuum polarization calculated using the QED operator model of Shabaev et al. (2018). These predictions demonstrate a reasonable agreement with our experimental data, with the largest discrepancy of about 100 meV observed for line 3C. We compared our results with other CI+MBPT data available in the literature (Wang et al., 2016; Santana et al., 2015; Gu, 2005), showing maximum deviations of up to 0.5 eV. The origin of the discrepancy between our CI+MBPT calculations and the previously published ones is unclear. We have also observed departures from the predictions of multiconfiguration Dirac-Fock (MCDF) and autostructure (AS) calculations (Aggarwal et al., 2003; Loch et al., 2005; Liang & Badnell, 2010), with deviations reaching up to 1–3 eV. However, we note that the atomic structures used in these calculations were necessarily small to facilitate their use in R-matrix collision calculations, which are computationally more demanding compared to atomic structure calculations. Other accurate predictions for Fe XVII from multireference Møller-Plesset (MRMP) are reported in Del Zanna & Ishikawa (2009) and included in the CHIANTI code. They show very good agreement with our experimental data.
@@
We compare our results with laboratory data from Beiersdorfer & Wargelin (1994) and Brown et al. (1998). Both works measured electron-impact spectra of Fe XVII under similar experimental conditions in the Lawrence Livermore National Laboratory (LLNL) EBIT facility using a crystal spectrometer employing a cesium acid phthalate (CsAP) crystal for the wavelength range of the lines presently discussed. Both measurements have carefully concatenated several spectra from different wavelength ranges and calibrated them against reference lines of hydrogenic and heliumlike oxygen, fluorine, and neon, similar to our work. Nevertheless, these two measurements are themselves marginally inconsistent with each other within their quoted uncertainties. Furthermore, we find that our measurements are also marginally inconsistent with both these previous measurements within uncertainties. The source of these marginal inconsistencies is unknown.
@@
We also compare our results with data from widely used databases and plasma codes. The NIST Atomic Spectroscopy Database (ASD) (Kramida et al., 2022) values showed significant deviations for lines 3C and 3A. However, when critically evaluated data by the authors of the NIST ASD (Kramida, 2019) were considered, we found a much better agreement with our experimental results (see Tab. 2). Comparison with AtomDB (Foster et al., 2012), CHIANTI (Del Zanna et al., 2021), and SPEX (Kaastra et al., 1996) databases and plasma codes revealed discrepancies as large as 1–2 eV. SPEX numbers showed better agreement with our results than those found in AtomDB, since SPEX has updated Fe-L atomic data (Gu et al., 2019, 2020, 2022), which were mainly calculated using FAC. Although the Astrophysical Plasma Emission Database (APED) version in AtomDB shows different values in its online webguide version (2.0.1) 222http://www.atomdb.org/Webguide/webguide.php and its pyatomdb version (3.0.4), the theoretical source is in both cases Loch et al. (2005), which uses the CI method, and disagrees by up to 1–2 eV from our results, as shown in Tab. 2. CHIANTI provides two sources for the energies: Autostructure (AS) theory (Liang & Badnell, 2010), and the more accurate set of data from MRMP theory (Del Zanna & Ishikawa, 2009). Note that the theoretical energy level data in AtomDB and Chianti are not used when generating model spectra when more accurate experimental or observational values exist. AtomDB replaces the most important transition energies of Fe XVII with the Brown et al. (1998) values, whereas CHIANTI uses the observed transition energies from solar Observation (Del Zanna & Ishikawa, 2009).
@@
Overall, most experimental and observational data agree with our experiment within 0.1 eV on average, well within the error bars of earlier works. However, there are substantial discrepancies with predictions from certain theoretical models, exceeding the margins of error associated with the experimental results. This highlights the urgent need to update the aforementioned databases to avoid pitfalls in astrophysical spectrum modeling and interpretation of observational data.
@@ 4 Summary and Conclusions
We presented high-precision transition-energy measurements of eight strong, astrophysically preeminent Fe XVII transitions required for plasma diagnostics. Our approach combined resonant photoexcitation of Fe XVII and narrow H-like and He-like transitions with high-resolution photoelectron spectroscopy (Rossnagel et al., 2001). This eliminates a very common source of systematic errors found even in advanced monochromators, namely quasiperiodic encoder interpolation errors (Follath & Balzer, 2010; Krempaský et al., 2011). As a result, our Fe XVII measurements represent a significant improvement in accuracy compared to previous experimental references, achieving an average enhancement of almost an order of magnitude. The uncertainties now stand at 10–15 meV, which translates to Doppler shifts of approximately km s-1. A further improvement in accuracy by another order of magnitude will require incorporating high-resolution/-stability voltage sources and more accurate voltmeters at ASPHERE to eliminate systematic errors associated with knowledge of the bias voltages.
@@
We have also improved our high-precision calculations by an order of magnitude in comparison with previous best calculations (Kühn et al., 2022; Cheung et al., 2021). This improvement allowed us, for the first time, to test the accuracy of QED corrections to the transition energies of a complicated 10 electron system. We expect that the achieved QED accuracy is applicable to a broad range of ions of intermediate degrees of ionization that can be treated with our large-scale CI or CI+all-order approaches. This has significant implications for predicting energy levels in systems where no experimental data are available for a wide range of applications in astrophysics, plasma physics, and atomic clock development (King et al., 2022). The established QED accuracy is deemed sufficient for high-precision prediction of HCI clock transitions (Kozlov et al., 2018). Improved accuracy of the experimental values would allow us to further decouple the uncertainty due to basis-set convergence from the uncertainty in the QED and improve theory predictions.
@@
Our improved transition-energy measurements for Fe XVII are sufficiently accurate that the uncertainties are no longer a significant part of the error budget for present or future planned astrophysical instruments, such as Chandra HETGS, XMM-Newton RGS, XRISM (Tashiro et al., 2018), Athena (Pajot et al., 2018), LEM (Kraft et al., 2022), HUBS (Cui et al., 2020), Arcus (Heilmann et al., 2022), HiReX (Nicastro et al., 2021), and Lynx (Schwartz et al., 2019). Future campaigns of similar measurements of prominent transitions in key ions (especially Fe-L shell ions) would be of great utility and could easily be directly included in commonly used astrophysical plasma spectral databases.
@@
The closeness of our large CI calculations to our measured values, with the worst deviation at line 3F of 33 meV amounting to a Doppler shift of only 13 km s-1, shows that such well-converged calculations are sufficiently accurate to be readily used in spectral databases. While there is no reason for this in the case of the lines measured in the present work, when accurate measurements are not available, similarly well-converged results could be used for other transitions of Fe XVII and many other ions. By including a very large number of configurations, our agreement becomes significantly better than that of other well-performing methods, such as results from less-converged large CI, MBPT, and MRMP calculations. For Fe XVII, our calculations are more accurate than even the best measurements for Fe-L shell transitions in Li-like through F-like ions (Brown et al., 2002). This suggests a near-future research program composed of comprehensive large CI calculations of transition energies for all ions of astrophysical interest up to Ne-like, supplemented by targeted experiments aimed at measuring the most important transition energies.
@@ Appendix A Large-scale CI calculations
The low-lying energies and wave functions are determined by solving the time-independent multielectron Schrödinger equation
Expanding the previous work (Kühn et al., 2022), we perform several calculations optimizing the basis-set convergence, including higher partial waves up to orbitals, and significantly expanding the set of reference configurations until convergence is reached in these parameters as well.
We have shown a comparison of our theoretical results for six transitions measured in this work with the experiment in the main text. We note that such a larger-scale computation of the 4C and 4D levels is beyond the capabilities of available computational resources (32 TB of memory and about 2000 CPUs on our largest available partition). Computing higher-lying levels requires computing of all the lower-lying levels with the same angular momentum and parity, drastically increasing memory requirement.
@@
We will discuss a complete assessment of the main contributions to the energies, including the basis-set construction, the inclusion of extra configurations, and QED. We find excellent agreement with experiments for all energies, at the level of 0.0004% for some levels. With the high level of accuracy attained, we are able to test QED contributions in the calculations of multielectron systems for the first time.
@@
The base calculation for the energy levels is done with a basis set and is listed in cm-1 in Tab. 4 in column “”. The contributions to the energy levels from expanding the basis set to and are in the columns “” and “”, respectively. The largest difference between the and calculations was 3 cm-1, so the basis set at the level of partial waves is considered sufficiently saturated. We note that although the basis was also used in Kühn et al. (2022), we constructed more compact basis in the present work, to significantly improve convergences with the principal quantum number . The basis is constructed in the 5 a.u. cavity, while the basis in Kühn et al. (2022) was constructed in a 20 a.u. cavity, with additional differences in the constructions of the higher partial-wave orbitals. A detailed comparison of the two computations confirms much better convergence properties of the present basis. We note very large computational resources needed for a basis-set expansion, especially for the inclusion of higher partial waves.
Contributions to higher partial waves are considered in the next six columns of Tab. 4. We calculated the contributions of extending the base basis set to include up to orbitals and listed them in column “”. Next, we successively increase the principal quantum number and increase the basis set up to . The contributions from orbitals and orbitals are given in columns “” and “”, respectively. The largest difference between and calculations was 9 cm-1, so the energies of including the higher orbitals have also converged sufficiently. We note that a large fraction of the contribution comes from very high- orbitals, so the inclusion of the first few orbitals does not give correct results for this partial wave. This effect is exacerbated for the and orbitals, where more of the contribution is expected to come from even with the present compact basis.
@@
We note that we have performed detailed convergence studies computing a separate contribution for each for the last few principal quantum numbers to evaluate convergence. Based on these data, we conservatively estimate the missing higher orbital contribution at 5 cm-1, higher orbital contribution at 20 cm-1, and higher orbital contribution at 50 cm-1. It appears that is not sufficiently converged. Table 4 shows that the contribution of all orbitals is about 1/2 of the contribution. Conservatively assuming a similar convergence pattern for higher partial waves gives 70 cm-1 for the partial wave and a similar total contribution for all the other partial waves. The total uncertainty due to the convergence of the basis set is then of the order of 230 cm-1. However, we note that the incomplete convergence of the basis is expected to cause a systematic shift of data for all levels; i.e., all energy values will be larger, with some smaller variances between the levels. It is possible that the partial-wave convergence is faster and the overall shift is smaller; therefore, we only use the above estimate to make an accuracy evaluation but do not shift the theory values. We note that overall +100 cm-1 shift of all of our values would improve the agreement of our data with the present experiment; however, this is the level of the experimental precision at 1 so improved experimental precision is needed to definitively test the basis set-convergence.
@@
Additionally, an extensive evaluation of the configuration weights was done to include important configurations in the list of basic reference configurations used to construct a final set of configurations. The weights of configurations signify the configuration’s contribution to the corresponding wave functions, and are calculated for each configuration as from Equation A1. These calculations are done by allowing single and double excitations to a much smaller basis set since the size of the computational problem will become prohibitive when additional reference configurations are included. The total contributions to including these extra configurations are given in the column “Extras” in Tab. 4. Beyond the initial 2 even- and 3 odd-parity configurations, we systematically included an additional 12 even- and 9 odd-parity reference configurations. Note that energies were calculated only for two even-parity levels to save computational resources and allow for additional reference configurations. The inclusion of these extra configurations contributes about 100 cm-1 shift to the energies and accounts for an additional 2 million relativistic configurations. Note also that these contributions would also be higher if the calculations were done with a larger basis set. We estimate an uncertainty from the convergence of the CI configuration set at the level of 50 cm-1, which is essentially negligible in comparison with the basis-set convergence uncertainty. We note that missing contributions can be both positive and negative in this case.
@@
Analysis of contributions to the 3F–3G, 3C–3D, and 3A–3B line differences given in the main text Table 3 shows that the basis-set expansion contribution effectively cancels for similar configurations; it is less than 50 cm-1 for all three cases. We also find that QED contributions play a major role in the 3F–3G energy difference. For 3C–3D, the contributions from the basis-set expansion and the addition of extra configurations essentially cancel out the QED. In the 3A–3B difference, the basis-set and extra configuration contributions cancel, leaving a shift from the QED. Therefore, comparing the differences in the energy values for similar configurations provides important additional information. It would be very useful to improve the uncertainty of the experiment as well as carry out such comparison in other ions with different degrees of ionization with 7–10 electrons.